OpenStudy (anonymous):

\Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = -1/2x + 4.

OpenStudy (anonymous):

y-4=2(x+2) y-4=2x+4 y=2x+8

OpenStudy (aroub):

y=mx+b , m= 2 , (-2,4) 4=2(-2)+b 4=-4+b 8=b the equation si y=2x+8